OpenStudy (anonymous):

Help Find all solutions in the interval [0, 2π). cos2x + 2 cos x + 1=0

OpenStudy (anonymous):

convert cos2x into cos^2x-sin^2

OpenStudy (anonymous):

now reformulate it without sin

OpenStudy (anonymous):

then use cos as X, then solve quadratic equation))

OpenStudy (anonymous):

Okay so would the ansewr be \[x=\frac{ \pi }{ 2 }, \frac{ 3\pi }{ 2}\]?

OpenStudy (anonymous):

is it \(\cos^2(x)\) or \(\cos(2x)\) for the first term?

OpenStudy (anonymous):

cos^2(x)

OpenStudy (anonymous):

then it is easier

OpenStudy (anonymous):

\[\cos^2(x)+2\cos(x)+1=\left(\cos(x)+1\right)^2\]

OpenStudy (anonymous):

therefore \(\cos(x)+1=0\iff \cos(x)=-1\) solve that one

OpenStudy (anonymous):

x=2pi?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(\cos(2\pi)=1\)